OpenStudy (anonymous):

f tanx+sinx=m,tanx-sinx=n show that m^2-n^2=4(mn)^0.5.

OpenStudy (anonymous):

whr r u stuck

OpenStudy (anonymous):

f tanx+sinx=m,tanx-sinx=n show that m^2-n^2=4(mn)^0.5.

OpenStudy (anonymous):

lol, plug in the values of m and n

hartnn (hartnn):

before plugging in write m^2-n^2 as (m+n)(m-n) it will be simplified.

hartnn (hartnn):

m+n=2 tan x m-n = 2 sin x

hartnn (hartnn):

so Left side = 4sin x tan x

hartnn (hartnn):

mn = tan^2 x - sin ^2 x using (a+b)(a-b)=a^2+b^2

hartnn (hartnn):

which is sin^2 x tan^x so \(\sqrt{mn}\)= sin x tan x

hartnn (hartnn):

which gives u what u wanted.